OpenStudy (anonymous):

What makes y = x2 + 3x + 2 different from y = x2 + 2x + 1

OpenStudy (anonymous):

the 3x and 2x and then the 2 and 1

OpenStudy (anonymous):

one # is a +2 and the other is a +1 then the other difference is that theres no 2 in one problem and theres no 3 in the other problem

OpenStudy (gw2011):

One way of looking at this is to factor both equations. y = x^2 + 3x + 2 = (x + 2) (x + 1) y = x^2 + 2x + 1 = (x + 1) (x + 1) = (x + 1)^2